OpenStudy (anonymous):

Fan and Medal If Cathy walks for 2 h and rides her bicycle for 1 h, she can travel 36 km. If she walks for 2 h and rides her bicycle for 2h, she can travel 56 km. How fast can she walk? How fast can she ride her bicycle? Let w represent walking and b represent biking. A. Write a system of equations. B. Solve the system. C. State the answers with the proper label or labels.

OpenStudy (tootsi123):

I would say A because you can't really solve this problem without an equation.

OpenStudy (anonymous):

2x+y=36 2x+2y=56

OpenStudy (tootsi123):

do you need to solve it to?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2x+y=36 2x+2y=56 Y=36-2x 2x+2(36-2x)=56 2x+72-4x=56 -2x=-16 X=8 Y=36-2(8) Y=20 So X=8 and Y=20

OpenStudy (anonymous):

She can walk 8 km fast and ride 20 km fast